A wooden toy kit has different painted solids. One solid is a cylinder 2cm and height 14 cm. what is the surface area of the cylinder and one can of paint covers 40m2. How many cylinders can be painted with one coat of paint.
Answer by amarjeeth123(569)

One solid is a cylinder 2cm and height 14 cm.
Surface area=2πrh+2πr^2=2π(2)(14)+2π(2)^2=56π+8π=64π square cm.=0.02010 square m
No of cylinders which can be painted=40/0.02010=1989.43
Answer=1989.43